# The adjacent Hindman’s theorem and the $\mathbb{Z}$-Ramsey theorem

### v1

@@ 3.1. The apartness condition
    Does -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$ (equivalently, $\mathnormal{\mathsf{AHT}}$) imply -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}_{s}$ (equivalently, $\mathnormal{\mathsf{AHT}}_{a}$)?
@@
    Can one prove, under $\mathnormal{\mathsf{RCA}_{0}}$ (or possibly under $\mathnormal{\mathsf{RCA}_{0}}+\mathsf{B}\Sigma_{2}^{0}$) that $\mathnormal{\mathsf{AHT}}$ (equivalently, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$) imply $\mathsf{D}^{2}$?
@@
The principle $\mathsf{D}^{2}$ mentioned in point (2) of Questions 15 is the statement, as considered in [3], that for every stable $c:[\mathbb{N}]^{2}\longrightarrow k$ (where “stable” means for each $n$, the function $m\longmapsto c(\{n,m\})$ is eventually constant) there exists an infinite set $X\subseteq\mathbb{N}$ and a colour $i<k$ such that for each $n\in X$, we have ${\lim_{m\to\infty}c(\{n,m\})=i}$; this statement is weaker than $\mathsf{IPT}^{2}\Rightarrow\mathsf{D}^{2}$ by a result of Dzhafarov and Hirst [3]. On the other hand, regarding point (3) of Questions 15, it is worth noting that D. Tavernelli [10, Section 2.3] has obtained results showing that the versions of $\mathnormal{\mathsf{AHT}}$ with more colours are not Weihrauch-reducible to the versions with fewer colours, although this still leaves open the possibility of having some version of $\mathnormal{\mathsf{AHT}}$ with more colours being provable (over $\mathnormal{\mathsf{RCA}_{0}}$, or maybe over $\mathnormal{\mathsf{RCA}_{0}}+\mathsf{B}\Sigma_{2}^{0}$) from another version with fewer colours.

### v2

@@ 3.1. The apartness condition
    Does -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$ (equivalently, $\mathnormal{\mathsf{AHT}}$) imply $\mathnormal{\mathrm{sep}\mathbb{Z}\text{-}\mathsf{RT}}^{2}$ (equivalently, $\mathnormal{\mathrm{ap}\mathsf{AHT}}$)?
@@
    Can one prove, under $\mathnormal{\mathsf{RCA}}_{0}$ (or possibly under $\mathnormal{\mathsf{RCA}}_{0}+\mathsf{B}\Sigma_{2}^{0}$) that $\mathnormal{\mathsf{AHT}}$ (equivalently, -$\mathnormal{\mathbb{Z}\text{-}\mathsf{RT}}^{2}$) implies $\mathsf{D}^{2}$?
@@
The principle $\mathsf{D}^{2}$ mentioned in point (2) of Questions 15 is the statement, as considered in [6], that for every stable $c:[\mathbb{N}]^{2}\longrightarrow k$ (where “stable” means for each $n$, the function $m\longmapsto c(\{n,m\})$ is eventually constant) there exists an infinite set $X\subseteq\mathbb{N}$ and a colour $i<k$ such that for each $n\in X$, we have ${\lim_{m\to\infty}c(\{n,m\})=i}$; this statement is implied by $\mathsf{IPT}^{2}$ by a result of Dzhafarov and Hirst [6]. On the other hand, regarding point (3) of Questions 15, it is worth noting that D. Tavernelli [17, Section 2.3] has obtained results showing that the versions of $\mathnormal{\mathsf{AHT}}$ with more colours are not Weihrauch-reducible to the versions with fewer colours, although this still leaves open the possibility of having some version of $\mathnormal{\mathsf{AHT}}$ with more colours being provable (over $\mathnormal{\mathsf{RCA}}_{0}$, or maybe over $\mathnormal{\mathsf{RCA}}_{0}+\mathsf{B}\Sigma_{2}^{0}$) from another version with fewer colours.